User name C. Fernandez-Ramirez
Log entry time 00:00:45 on December 6,2007
Entry number 218443
keyword=Swing shift summary
16:00 - Work at Hall finished. Beam permit
16:00 - Beam requested, they cannot deliver beam yet. Waiting...
17:00 - Left spectrometer cycled
18:00 - Beam on
18:15 - BPMB -> X=-1.9 : Y= 0.02
18:35 - Harp scan -> 0.06x0.065
20:30 - Harp scan -> 0.175x0.152
20:40 - Problems with the compton polarimeter. MCC trying to fix it
21:05 - Problems in the accelerator. Beam off and off and off...